MATH 574, NOTES 12 FINAL REVIEW

Things Not to Know:

▶ Example (3), Notes 1 ▶ Example (6), Notes 1 ▶ Homework (5), Notes 1 ▶ Example (7), Notes 2 ▶ Example (8), Notes 2 ▶ Example (5), Notes 4 ▶ Example (6), Notes 4 ▶ Example (7), Notes 4 ▶ Example (12), Notes 5 ▶ Example (7), Notes 7 ▶ Proof in Example (6), Notes 9 ▶ Proof in Example (7), Notes 9 ▶ Proof in Example (8), Notes 9 ▶ Example (2), Notes 10 ▶ Example (4), Notes 11

Things to Know:

▶ All remaining past problems - examples

- homework

- quizzes

- tests

- practice problems

Comment: At least 50% of the final will be past problems (with possible number variations).